List the different types and formulas for the following ratios:
2.1 Capital Structure (Leverage) Ratios (10)
2.2 Profitability Ratios (10)
2.3 Liquidity Ratios
1
Expert's answer
2021-09-15T17:16:01-0400

2.1 Capital Structure (Leverage) Ratios

1.Debt-to-Equity Ratio"=\frac{Total Debt}{Total Equity}"


2.Debt-to-Assets Ratio"=\frac{Total Debts}{Total Assets}"


3.Debt-to-Capital Ratio"=\frac{Today Debt}{Total Debts + Total Equity}"


4.Assets-to-Equity Ratio"=\frac{Total Assets} {Total Equity}"


5.Debt-to-EBITDA Ratio"=\frac{Total Debt}{Earnings Before Interest Taxes Depreciation and Amortization}"


2.2 Profitability Ratios


1.Return on Assets"=\frac{Net Income} {Average Total Assets}"

2.Return on Equity"=\frac{Net Income} {Average Stockholder Equity}"


3.Profit Margin"=\frac{Net Income} {Sales}"


4.Earnings Per Share"=\frac{Net Income} {Number of Common Share Outstanding}"


2.3 Liquidity Ratios

1.Current Ratio"=\frac{Current Assets} {Current Liabilities}"


2.Quick Ratio"=\frac{Quick Assets} {Current Liabilities}"


3.Net Working Capital Ratio"=\frac{Current Assets - Current Liabilities}{Total Assets}"


4.Cash Ratio"=\frac{Cash}{Current Liabilities}"


5.Operating Cash Flow Ratio"=\frac{Operating Cash Flow} {Current Liabilities}"


6.Cash Coverage Ratio"\frac{Earnings Before Interest and Taxes + Depreciation} {Interest}"